# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v1

@@ 2.3 EMA Teacher Network
In SRA, we do not need a off-the-shelf teacher to give the prior guidance. Meanwhile, using the output of the same model as the target to conduct supervision would cause training instability [11, 25]. Thus, we follow some works in self-supervised learning (SSL) [78, 8, 27] that build the teacher from past iterations of the student network using an exponential moving average (EMA) on the student weights. In specific, the update role of EMA is $\theta_{t}=\alpha\theta_{t}+(1-\alpha)\theta_{s}$, where $\alpha\in[0,1)$ is the momentum coefficient. We study different values of $\alpha$ in Section 3.2 and show that $\alpha=0.9999$ unchanged works surprisingly well in our framework. Moreover, we do not use other operations like clustering constraints [6, 7], batch normalizations [25, 61], and centering [8, 78] in SSL, because we find that the training progress is already stable enough without applying these tricks.
@@ 3 Experiment
  • Dose SRA work on different baselines across different model size? (Figure 4)
  • Can SRA show comparable or superior performance against other methods that leverage either representation training paradigm or representation foundation model? (Table 3.3)
@@ 3.1 Experimental Setup
Implementation details. Unless otherwise specified, the training details are strictly followed the setup in DiT [57] and SiT [52], including AdamW [51] with a constant learning rate of 1e-4, no weight decay, batchsize of 256, using the Stable Diffusion VAE [62] to extract the latent, and etc. We use ImageNet [16], where each image is preprocessed to the resolution of 256$\times$256 (denoted as ImageNet 256$\times$256), and follow ADM [17] for other data preprocessing protocols. For model configurations, we use the B/2, L/2, and XL/2 architectures introduced in the DiT and SiT papers, which process inputs with a patch size of 2. Additional experimental details, hyperparameter settings, linear probing details, and computing resources, are provided in Appendix 3.
Evaluation. We report Fréchet inception distance (FID [28]), sFID [54], inception score (IS [64]), precision (Pre.) and recall (Rec.) [41] . To ensure a fair comparison with previous methods, we also use the ADM’s TensorFlow evaluation suite  [17]with 50K samples and the same reference statistics. More details of each metric are provided in Appendix D.
Sampler. For DiT [57], we use DDPM sampler and set the number of function evaluations (NFE) as 250 by default. For SiT [52], we use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the NEF as 250 by default. The settings also align with those used in DiT and SiT papers.
Baselines. We use several recent diffusion-based generation methods as baselines. Specifically, we consider the following three types of approaches: (a) Pixel diffusion: ADM [17], VDM+++++ + [37], Simple diffusion [32], CDM [30], (b) Latent diffusion with U-Net: LDM [63], and (c) Latent diffusion with transformers: DiT [57], SiT [52], SD-DiT [79], MaskDiT [77], TREAD [39], REPA [74], and MAETok [9]. Here in diffusion transformers family, we choose to compare with DiT/SiT and their modifications with representation components involved but do not compare with works aiming at designing advanced architecture like lightningDiT [73] and DDT [67]. Detailed descriptions of each baseline method are provided in Appendix E.
@@ 3.2 Component-Wise Analysis
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $0\sim k_{max}$ denotes time Interval $k$ is a random value in $[0,k_{max})$. PH. denotes whether to use projection head.
@@
Block layers for alignment. We begin by analyzing the effect of using different blocks of student and teacher for alignment. We observe that using the teacher’s last but not least few layers (e.g., 8) to regulate the student’s first layers (e.g., 3) leads to optimal performance. We assume that the first few layers need more guidance so they can catch semantically meaningful representation for subsequent generation. Meanwhile, there is a strong correlation between the quality of the representations of the teacher’s layers and the performance of the corresponding aligned student (we give the quantitative results and analysis in latter Section 3.4). Based on these results, we set alignment layers as $83\xrightarrow{}$ 8, $166\xrightarrow{}$ 16, and $208\xrightarrow{}$ 20 for B, L, and XL models respectively as defaultFor DiTs, we set alignment layers as $73\xrightarrow{}$ 7, $146\xrightarrow{}$ 14, and $168\xrightarrow{}$ 16 by default because DiT’s discriminative behavior across layers is sightly different from SiT’s (see Figure 2 and Figure 7).
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here we study fixed and dynamic interval. Notably, we find that using teacher’s features input with lower noise than student’s leads to performance enhancement and interval value with 0.1 or the mean with 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance; but an excessively large time interval can hinder the model’s learning process, causing it to only focus on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor 0\sim 200\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
@@
Teacher network for alignment. In other generative learning studies, the EMA model is often employed solely for evaluation. However, as we need it to provide guidance during training, we study different updating methods here. Here, we investigate several different strategies to build the teacher. First, we find that using teacher copied from student would impair the performance, which testifies our argument in Section 1. Next, we consider using strategy used in self-supervised learning works [8, 78, 25] that updates momentum coefficient from 0.996 to 1 during training. However, in our framework, this does not work well. Finally, we use momentum coefficient of 0.9999 unchanged which is commonly used in other generative learning works [17, 57, 52] and find it is also suitable for our framework. Thus, we set momentum coefficient as 0.9999 in future experiments.
Effect of projection head for alignment. We finally examine the effect of the projection head for alignment. Surprisingly, We observe that using this simple head to post-possess the student’s output is much better than directly using it to alignment. We hypothesize this slight operation enables the model to learn more effective hidden representations for subsequent projection head to conduct transformation for final alignment, rather than explicitly aligning the entire latent feature that could potentially disrupt the original generation field that each layer and timestep is responsible for [76, 75, 22]. Hence, we keep using the projection head in future experiments.
@@ 3.3 System-Level Comparison
Based on the analysis, we perform a system-level comparison between recent state-of-the-art diffusion model approaches and diffusion transformers with SRA.
First, we compare the FID values between vanilla DiT or SiT and the same models trained with SRA. As shown in Figure. 4, diffusion transformers trained with SRA demonstrate substantial improvements in performance at each training step across different types , as well as various sizes. Moreover, similar to the observation in some SSL works [56, 21], we notice that the effect of SRA in larger size model is more significant, which probably because larger model tends to provide richer guidance. It is also worth noting that the benefits of SRA do not saturate even when the models have already achieved a low FID score. We assume that this is likely due to the teacher’s constantly improving capacity throughout training, which allows it to provide better and better representation guidance for the student when the training goes on.
Figure 4: FID comparisons with vanilla DiTs and SiTs across different model size on ImageNet 256$\times$256 without classifier-free guidance (CFG).
@@
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent dif- fusion model methods using classifier-free guidance [31]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior than methods (e.g., MaskDiT [77]) that rely on auxiliary and complex representation training frameworks, and is comparable with methods (e.g., REPA [74]) that depend heavily on powerful external representation priors. Moreover, due to the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
@@ 3.4 Ablation Study
In this part, we aim at testing whether representation truly matters in SRA. We give the answer by conducting following experiments.
@@
Tight coupling between generation quality and representation guidance in SRA. We then investigate the correlation between generation performance and the representation guidance (detailed experimental setup can be find in Appendix G) in SRA. Figure 6c reveals a strong correlation between linear probing accuracy and FID scores as the teacher network layers for alignment are varied. This finding underscores that the model’s generative capabilities are indeed closely tied to the effectiveness of the self-representation guidance mechanism.
@@ 4 Related Work
We discuss with the most relevant studies here and provide a more discussion in Appendix I.
Representations guidance for diffusion transformer’s pre-training. Many recent works have attempted to introducing representation guidance in diffusion transformer’s training. MaskDiT [77] and SD-DiT [79] add MAE’s [26] and IBOT’s [78] training paradigm into the original DiT’s training progress. TREAD [39] design a token routing strategy with MAE loss to speed-up diffusion model’s training. REPA [74] utilize a large-scale data pre-trained representation model to regulate diffusion model’s latent feature. VA-VAE [73] and MAETok [9] align the latent distribution of the tokenizer with additional representation foundation model and show this alignment is beneficial to resulting diffusion models. Different from these work, we aim at looking for representation guidance in the diffusion model itself and it’s own training paradigm.
EMA model as teacher guidance. Different from traditional knowledge distillation [24] that uses a stronger pre-trained model as teacher, using EMA model [33] as teacher can be seen as self-distillation because the weight of the teacher model is obtained by weighted moving average of the parameters of the student model. This method needs a feasible pretext task to succeed. For example, MoCo [27, 13] sets the EMA teacher as momentum encoder and use contrastive learning to guidance the student model; DINO [8, 56] feeds two views of images to EMA teacher and trainable student then forces the student’s output distribution to be close to teacher’s. Our work also shares some similarities, where we set aligning student model’s latent feature in earlier layer and higher noise with that in later layer and lower noise of the EMA teacher as our pretext task for training.
@@ 5 Conclusion and Discussion
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns the output latent representation of the diffusion transformer in earlier layer with higher noise to that in later layer with lower noise to progressively enhance the overall representation learning during only generative process.
@@ Appendix B Descriptions for Two Types of Baseline Models
In this paper, we use DiT and SiT as our baseline models. We now provide an overview of two types of generative models that are variants of denoising autoencoders and are used to learn the target distribution. Specifically, we discuss Denoising Diffusion Probabilistic Models (DDPMs) like DiT in Section B.1 and Stochastic Interpolants Models like SiT in Section B.2.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise diffusion-based models [29, 55] aim to model the target distribution ( $p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
For a given ${\mathbf{x}}_{0}\sim p({\mathbf{x}})$, the forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})$ ) is defined as: [ $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$, ] where $\beta_{t}\in(0,1)$ are pre-defined small hyperparameters. DDPM [29] formalizes the reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) as:
$p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})=\mathcal{N}\Big{(}{\mathbf{x}}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\big{(}{\mathbf{x}}_{t}-\frac{\sigma_{t}^{2}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\bm{\theta}}({\mathbf{x}}_{t},t)\big{)},\mathbf{\Sigma}_{\theta}({\mathbf{x}}_{t},t)\Big{)},$ (8)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$, and $\bm{\epsilon}{\bm{\theta}}({\mathbf{x}}_{t},t)$ is parameterized by a neural network.
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{\ast},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (9)
For the covariance term ( $\mathbf{\Sigma}{\bm{\theta}}({\mathbf{x}}_{t},t)$ ), DDPM[29] demonstrated that setting it as ( $\sigma_{t}^{2}\mathbf{I}$ ) with ( $\beta_{t}=\sigma_{t}^{2}$ ) is sufficient. Subsequently, Improved-DDPM[55] showed that performance can be enhanced by jointly learning ( $\mathbf{\Sigma}{\bm{\theta}}({\mathbf{x}}_{t},t)$ ) along with ( $\bm{\epsilon}{\bm{\theta}}({\mathbf{x}}_{t},t)$ ) in a dimension-wise manner using the following objective:

### v3

@@ 2.3 EMA Teacher Network
In SRA, we do not need an off-the-shelf teacher to give the prior guidance. Meanwhile, using the output of the same model as the target to conduct supervision would cause training instability [12, 28]. Thus, we follow some works in self-supervised learning (SSL) [85, 8, 30] that build the teacher from past iterations of the student network using an exponential moving average (EMA) on the student weights. In specific, the updated role of EMA is $\zeta_{t}=\alpha\zeta_{t}+(1-\alpha)\zeta_{s}$, where $\alpha\in[0,1)$ is the momentum coefficient. We study different values of $\alpha$ in Section 3.2 and show that $\alpha=0.9999$ unchanged works surprisingly well in our framework. Moreover, we do not use other operations like clustering constraints [6, 7], batch normalizations [28, 67], and centering [8, 85] in SSL because we find that the training progress is already stable enough without applying these tricks.
@@ 3 Experiment
  • Dose SRA work on different baselines across different model sizes? (Figure 4)
  • Can SRA show comparable or superior performance against other methods that leverage either the representation training paradigm or the representation foundation model? (Table 3.3)
@@ 3.1 Experimental Setup
Implementation details. Unless otherwise specified, the training details strictly follow the setup in DiT [62] and SiT [56], including AdamW [54] with a constant learning rate of 1e-4, no weight decay, batchsize of 256, using the Stable Diffusion VAE [68] to extract the latent, and etc. We use ImageNet [18], where each image is preprocessed to the resolution of 256$\times$256 (denoted as ImageNet 256$\times$256), and follow ADM [19] for other data preprocessing protocols. For model configurations, we use the B/2, L/2, and XL/2 architectures introduced in the DiT and SiT papers, which process inputs with a patch size of 2. Additional experimental details, hyperparameter settings, linear probing details, and computing resources, are provided in Appendix 3.
Evaluation. We report Fréchet inception distance (FID [31]), sFID [59], inception score (IS [70]), precision (Pre.) and recall (Rec.) [44]. To ensure a fair comparison with previous methods, we also use the ADM’s TensorFlow evaluation suite  [19]with 50K samples and the same reference statistics. A detailed breakdown of each evaluation metric is included in Appendix D.
Sampler. For DiT [62], we use the DDPM sampler and set the number of function evaluations (NFE) as 250 by default. For SiT [56], we use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the NEF as 250 by default. The settings also align with those used in DiT and SiT.
Methods for Comparison. To compare performance, we utilize recent advanced methods based on diffusion models. In particular, We focus on three distinct categories of approaches: (a) Pixel diffusion: ADM [19], VDM+++++ + [40], Simple diffusion [35], CDM [33], (b) Latent diffusion with U-Net: LDM [69], and (c) Latent diffusion with transformers: DiT [62], SiT [56], SD-DiT [86], MaskDiT [84], TREAD [42], REPA [81], and MAETok [9]. Here in the diffusion transformers family, we choose to compare with DiT/SiT and their modifications with representation components involved but do not compare with works aiming at designing advanced architecture like lightningDiT [80] and DDT [74]. Detailed descriptions of each method can be found in Appendix E.
@@ 3.2 Component-Wise Analysis
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $0\sim k_{max}$ denotes time Interval $k$ is a random value in $[0,k_{max})$. PH. denotes whether to use the projection head.
@@
Block layers for alignment. We begin by analyzing the effect of using different blocks of student and teacher for alignment. We observe that using the teacher’s last but not least few layers (e.g., 8) to regulate the student’s first layers (e.g., 3) leads to optimal performance. We assume that the first few layers need more guidance so they can catch semantically meaningful representation for subsequent generation. Meanwhile, there is a strong correlation between the quality of the representations of the teacher’s layers and the performance of the corresponding aligned student (we give the quantitative results and analysis in the latter Section 3.4). Based on these results, we set alignment layers as $83\xrightarrow{}$ 8, $166\xrightarrow{}$ 16, and $208\xrightarrow{}$ 20 for B, L, and XL models respectively as defaultFor DiTs, we set alignment layers as $73\xrightarrow{}$ 7, $146\xrightarrow{}$ 14, and $168\xrightarrow{}$ 16 by default because DiT’s discriminative behavior across layers is slightly different from SiT’s (see Figure 2 and Figure 7).
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with lower noise than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor 0\sim 200\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
@@
Teacher network for alignment. In other generative learning studies, the EMA model is often employed solely for evaluation. However, as we need it to provide guidance during training, we study different updating methods. Here, we investigate several different strategies to build the teacher. First, we find that using teacher copied from a student would impair the performance, which testifies to our argument in Section 1. Next, we consider using the strategy used in self-supervised learning works [8, 85, 28] that updates the momentum coefficient from 0.996 to 1 during training. However, in our framework, this does not work well. Finally, we use momentum coefficient of 0.9999 unchanged, which is commonly used in other generative learning works [19, 62, 56] and find it is also suitable for our framework. Thus, we set the momentum coefficient as 0.9999 in future experiments.
Effect of projection head for alignment. We finally examine the effect of the projection head for alignment. Surprisingly, We observe that using this simple head to post-possess the student’s output is much better than directly using it to align. We hypothesize this slight operation enables the model to learn more effective hidden representations for subsequent projection head to conduct transformation for final alignment, rather than explicitly aligning the entire latent feature that could potentially disrupt the original generation field that each layer and timestep is responsible for [83, 82, 25]. Hence, we keep using the projection head in future experiments.
@@ 3.3 System-Level Comparison
In this section, a system-level comparative study is performed to assess recent denoise-based approaches against diffusion transformers with SRA.
First, we compare the FID values between vanilla DiT or SiT and the same models trained with SRA. As shown in Figure. 4, diffusion transformers trained with SRA demonstrate substantial improvements in performance at each training step across different types , as well as various sizes. Moreover, similar to the observation in some SSL works [61, 23], we notice that the effect of SRA in a larger size model is more significant, which is probably because the larger model tends to provide richer guidance. It is also worth noting that the benefits of SRA do not saturate even when the models have already achieved a low FID score. We assume that this is likely due to the teacher’s constantly improving capacity throughout training, which allows it to provide better and better representation guidance for the student when the training goes on.
Figure 4: FID comparisons with vanilla DiTs and SiTs across different model sizes on ImageNet 256$\times$256 without classifier-free guidance (CFG).
@@
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent dif- fusion model methods using classifier-free guidance [34]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT [84]) that rely on auxiliary and complex representation training frameworks and is comparable with methods (e.g., REPA [81]) that depend heavily on powerful external representation priors. Moreover, because of the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
@@ 3.4 Ablation Study
Since SRA introduces representation guidance in an implicit way, we thus aim at testing whether representation truly matters in SRA. We give the answer with the following experiments.
@@
Tight coupling between generation quality and representation guidance in SRA. We then investigate the correlation between generation performance and the representation guidance (detailed experimental setup can be found in Appendix G) in SRA. Figure 6c reveals a strong correlation between linear probing accuracy and FID scores as the teacher network layers for alignment are varied. This finding underscores that the model’s generative capabilities are indeed closely tied to the effectiveness of the self-representation guidance mechanism.
@@ 4 Related Work
Here, we highlight key related studies and defer a discussion of other relevant studies to Appendix I.
Representations guidance for diffusion transformer’s pre-training. Many recent works have attempted to introduce representation guidance in diffusion transformer training. MaskDiT [84] and SD-DiT [86] add MAE’s [29] and IBOT’s [85] training paradigm into the original DiT’s training progress. TREAD [42] design a token routing strategy with MAE loss to speed up the diffusion model’s training. REPA [81] utilizes a large-scale data pre-trained representation model for regulating the diffusion model’s latent feature. VA-VAE [80] and MAETok [9] align the latent distribution of the tokenizer with the external representation foundation model and show this alignment is beneficial to resulting diffusion models. Different from this work, we aim to look for representation guidance in the diffusion model itself and its own training paradigm.
EMA model as teacher. Unlike traditional knowledge distillation [27] that uses a stronger pre-trained model as the teacher, using EMA model [36] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [30, 14] sets the EMA teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [8, 61] feeds two views of images to the EMA teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer and higher noise with that in the later layer and lower noise of the EMA teacher as our pretext task for training.
@@ 5 Conclusion and Discussion
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns the latent representation of the diffusion transformer in the earlier layer with higher noise to that in the later layer with lower noise to progressively enhance the overall representation learning without external components.
@@
6 Acknowledgment
Our SRA manuscript referred to some of the color schemes and structural options provided in the REPA [81] manuscript. We sincerely thank REPA’s authors for their fully open-source project. We also thank Sihyun Yu, Sizhe Dang, and Zanyi Wang for the helpful discussions and suggestions during the progress of SRA project.
@@ Appendix B Descriptions for Two Types of Baseline Models
In this paper, we use DiT and SiT as our baseline models. We now provide an overview of two types of generative models that are variants of denoising autoencoders and are used to learn the target distribution. Specifically, we discuss Denoising Diffusion Probabilistic Models (DDPMs) like DiT in Section B.1 and Stochastic interpolant models like SiT in Section B.2.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise denoise-based models [32, 60] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
For a given ${\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0})$, the forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})$ ) is defined as: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$, where $\beta_{t}\in(0,1)$ are pre-defined small hyperparameters. DDPM [32] formalizes the reverse process $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ as:
$p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})=\mathcal{N}\Big{(}{\mathbf{x}}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\big{(}{\mathbf{x}}_{t}-\frac{\sigma_{t}^{2}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)\big{)},\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)\Big{)},$ (7)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$, and $\bm{\epsilon}_{\bm{\zeta}}(\bm{x_{t}},t)$ is parameterized by a neural network.
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (8)
For the covariance term ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ), DDPM[32] demonstrated that setting it as ( $\sigma_{t}^{2}\mathbf{I}$ ) with ( $\beta_{t}=\sigma_{t}^{2}$ ) is sufficient. Subsequently, Improved-DDPM[60] showed that performance can be enhanced by jointly learning ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) along with ( $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) in a dimension-wise manner using the following objective: